### v1

# Parity violation in atoms and electron scattering revisited

Abstract
Exchange by two neutrinos (as well as by other fermions) generates a long-range parity-violating (PV) potential of the form $\sim G^{2}/r^{5}$, with characteristic range $\hbar/(2m_{\nu}c)$. This interaction produces a $0.8\%$ correction to the effective weak charge of the Cs atom, thereby resolving the $2\sigma$ discrepancy between the Standard Model prediction and the Cs PV measurement, and yielding $\sin^{2}\theta_{W}=0.2375(19)$ at $q^{2}\approx 0$ (to be compared with the Standard Model value $\sin^{2}\theta_{W}=0.2387$). We derive updated limits on an extra $Z^{\prime}$ boson and obtain a constraint on isospin-conserving oblique radiative corrections characterized by the Peskin–Takeuchi parameter, S=0.32(53)S=-0.32(53) at $q^{2}\approx 0$. The corresponding relative correction to the proton weak charge is $3\%$, and a similar correction is expected for the electron weak charge.
I Introduction
Exchange by two neutrinos as in Fig. 1 produces long range potential $\sim G^{2}/r^{5}$ in the particle interaction, where $G$ is Fermi constant  [19, 18, 17, 25]. However, effects of parity conserving part of this potential are many orders of magnitude smaller than the sensitivity of the experiments [27, 1, 7, 46, 43, 41, 40, 32].
Figure 1: Feynman diagrams contributing to the long-range force due to neutrino exchange in atoms. Diagram (a) may also contain other fermions and quarks in the loop while the diagram (b) has only the electronic neutrino in the loop. Diagram (c) has been actually calculated at the quark level.
In Refs. [21, 24, 12, 23, 22] it was noted that the neutrino exchange potential has parity violating (PV) part $\sim G^{2}/r^{5}$ with the range $\hbar/2m_{\nu}c$ exceeding atomic size by several orders of magnitude. It was also demonstrated that mixed $Z-\gamma$ electron vacuum polarization produces PV potential $\sim\alpha G/r^{3}$ with the range $\hbar/2m_{e}c$=193 fm, which exceeds the range of the weak interaction by five orders of magnitude [21, 12], similar to the range of the Uehling potential due to electron vacuum polarization by the nuclear Coulomb field.
In the present paper we show that the contributions of the diagrams in Fig. 1 with exchange by two neutrinos or by other fermions to the PV effects in atoms is $\sim$1%, and this significantly exceeds the error 0.35% of the PNC measurement in Cs atom [49] and the error <<0.5% in the many-body atomic calculations of the $Z$-boson contribution [14, 10, 3, 4, 20, 34, 35, 9]. The work is in progress to improve both experimental [8, 44, 36] and theoretical [42] accuracy.
The error in atomic calculations cancels out in the ratio of the PV amplitudes in different isotopes of the same atom [11, 6, 48]. The measurement for the chain of isotopes of Yb atom was performed in Ref. [2].
The study of PV in atoms provides high precision tests of the Standard Model and new physics beyond it [37, 38].
We will also show that the diagrams on Fig. 1 give significant contributions to the PV effects in low-energy scattering of electron from proton, deuteron and other nuclei. The corresponding corrections to the proton and nuclear weak charges are found. Similar correction to PV effects in electron-electron (Møller) scattering and electron weak charge is expected.
Weak charge $Q_{W}$ is the constant of the electron-nucleon weak interaction due to the $Z$-boson exchange which has interaction range $r_{Z}=\hbar/M_{Z}c\approx 0.002$ fm. On the nuclear and atomic scales this may be considered as a Fermi-type contact interaction.
Radiative corrections to the proton and neutron weak charges of the order $\alpha\approx 1/137$ have been calculated in Refs. [29, 28, 30, 5] and were shown to exceed 1%. However, diagrams on Fig. 1, which actually give comparable contributions, have not been included into these calculations. The point is that at large distance these diagrams give interaction proportional to $G^{2}$, which is very small. However, this interaction is highly singular, $\sim G^{2}/r^{5}$, and integration over distance with electron wave functions and effective cut-off at $r\sim 1/M_{Z}$ gives contribution $\sim G^{2}M_{Z}^{2}\sim\alpha G$. Thus, the suppression relative to the ordinary weak interaction mediated by the $Z$-boson exchange is $\sim\alpha$ [12].
For electron-electron scattering there are also box diagrams with crossing electron and crossing neutrinos lines. Note that in the case of Majorana neutrinos, there is an additional box diagram with neutrinos as the vertical sides. Potentially, this could give a possibility to identify Majorana neutrinos using PV in electron scattering, but the contribution of this diagram is small.
To perform accurate calculations, we require the effective potential generated by the diagrams shown in Fig. 1 at distances $r\sim 1/M_{Z}$. The neutrino-exchange contribution to this potential, valid for all distances satisfying $r\ll 1/m_{\nu}$, was calculated in Refs. [23, 22], and our analysis is based on these results. We further include the additional contributions arising from other fermions in the loop.
II Effective potential due to the fermion loop
Parity violating interactions produce PV effects in atoms by mixing $s_{1/2}$ and $p_{1/2}$ electron orbitals. To find corrections produced by diagrams on Fig. 1 it is sufficient to calculate the ratio of the matrix elements
$\frac{\langle ns_{1/2}|W_{f}(r)|n^{\prime}p_{1/2}\rangle}{\langle ns_{1/2}|W_{Z}(r)|n^{\prime}p_{1/2}\rangle}\,$, (1)
where $W_{Z}$ and $W_{f}$ are the PV potentials produced by $Z$ boson tree-level exchenge and $f$ fermion loop correspondingly.
Dominating contribution to both matrix elements comes from the distance between electron and quark of order $r\sim 1/M_{Z}$, which is very small as compared to the scale of the variation of electron wave functions and nuclear size. Therefore, we may replace both potentials by their Fermi-type contact approximations $w_{Z,f}\delta(r)$ using relation $\int W_{Z,f}(r)4\pi r^{2}dr=w_{Z,f}\int\delta(r)d^{3}r$. Then the nuclear-spin-independent (NSI) part of the electron-nucleus PV operator $W(r)$ may be presented as
$W(r)=-\frac{G}{2\sqrt{2}}(Q_{W}+\delta Q_{W})\gamma_{5}\rho(r)\,$,\, (2)
where $Q_{W}=-N+(1-4\sin^{2}\theta_{W})Z$ is the weak nuclear charge, $\delta Q_{W}$ is the correction produced by the diagrams on Fig. 1, $\rho(r)$ is the nuclear density normalized by the condition $\int\rho(r)dV=1$, $N$ and $Z$ are numbers of neutrons and protons in the nucleus. Inclusion of radiative corrections from Refs. [29, 28, 30, 5] (see also [33, 16]) gives $Q_{W}=-0.98897N+(1.0250-3.9978\sin^{2}\theta_{W})Z$, which gives for the cesium atom the standard model value of the weak charge $Q_{W}^{SM}=-73.26(1)$, using $\sin^{2}\theta_{W}=0.23873$ for momentum $q^{2}\approx 0$ [16]. The experimental measurements of the parity violation in cesium atom give $\sin^{2}\theta_{W}=0.2349(18)$ [16].
Note that the differences in the experimental values of Cs weak charge and Weinberg angle in 2024 and 2022 editions of Particle Data group reviews are due to the different values of the Stark polarizability $\beta$ used (Cs PV experiment measures ratio of the PV amplitude to $\beta$). The 2023 paper [45] corrected the result of Ref. [47] and confirmed the original value of $\beta$ found in our paper [15].
Calculations of the neutrino exchange potentials, presented in Appendix, are similar to that described in detail in Ref. [23]. Let us start from the diagram (a) in Fig. 1. The interaction constant on the bottom fermion line cancels with the single $Z$-boson exchange contribution in the ratio (1). Therefore, for any fermion on the bottom line (e.g. proton, neutron or electron in the electron- electron scattering case) the ratio of first diagram to the $Z$-boson contribution is the same. As a result we obtain
$\frac{\delta Q_{W}^{(1)}}{Q_{W}}=-\frac{g^{2}}{96\pi^{2}c_{W}^{2}}N_{\text{eff}}=-0.86\%\,$, (3)
where $g\approx 0.65$ is the $Z$-boson-neutrino interaction constant, $c_{W}^{2}=1-s^{2}_{W}=0.761$, $s_{W}^{2}\equiv\sin^{2}\theta_{W}$. Following Refs. [41, 12], we introduced here effective number of neutrinos $N_{\text{eff}}=14.5$. Indeed, there are 3 ordinary neutrinos. However, we should add to the loop all fermions with mass smaller than $Z$-boson mass, including $e,\,\mu,\,\tau,\,u,\,d,\,s,\,c,\,b$. $N_{\text{eff}}=14.5$ is not an integer number since interaction constants of other fermions with $Z$-boson are different from that for neutrinos.
Diagrams (b) and (c) in Fig. 1 do not have enhancement factor $N_{\text{eff}}$. Using the calculations of the momentum and radial integrals presented in Appendix with the result in Eq. (31), we obtain the following correction to the nuclear weak charge:
$\frac{\delta Q_{W}^{(2,3)}}{Q_{W}}=1.7\times 10^{-3}\frac{N-Z}{N-(1-4s^{2}_{W})Z}\,$, (4)
For the proton weak charge this ratio, enhanced by the small value of $(1-4\sin^{2}\theta_{W})=0.044$ in the denominator, is 3.8%. Adding contribution (3) corresponding to the diagram (a), we obtain the total correction to the weak charge
$\frac{\delta Q_{W}}{Q_{W}}=\left(-0.86+0.017\frac{N-Z}{N-(1-4s^{2}_{W})Z}\right)\%\,.$ (5)
Equation (5) shows that the contribution of the diagrams (b) and (c) (given in the last term) for all atoms, except for hydrogen, is small, so we obtain $\delta Q_{W}/Q_{W}\approx-0.8\%$ in all medium and heavy atoms. For the proton weak charge this gives,
$\frac{\delta q_{W}^{p}}{q_{W}^{p}}\approx 3.0\%\,.$ (6)
Similar relative correction to the PV effect in muonium atom was obtained in Refs.  [23, 22], where the authors do not interpret this as a correction to the weak charge. The enhancement due to the small value of $(1-4s^{2}_{W})$ in denominator is also expected for the electron weak charge in the case of PV in electron-electron Møller scattering.
III Comparison with experimental data on weak charges and limits on new physics
If we use the value of the proton weak charge obtained in the recent very accurate lattice calculation, $Q_{W}^{SM}=0.06987(50)$, we obtain corrected value $Q_{W}^{SM}+\delta Q_{W}^{SM}=0.06987(1+0.03)=0.0719$. It agrees with $Q_{W}^{\text{exp}}=0.0719(45)$ measured by Qweak collaboration [26].
The correction to the electron weak charge is also expected at the level of several per cent. This does not contradict to the experimental value of the electron weak charge $Q_{W}^{\text{exp}}=-0.053(11)$ obtained by E158 collaboration [39]. The accuracy in the experimental measurement of the proton weak charge in the electron-proton scattering by the Qweak collaboration is 7%, the accuracy in the measurement of the PV asymmetry in electron-electron scattering by E158 collaboration is 20%. However, both experimental collaborations are expecting significant improvements of the accuracy soon.
The most sensitive point to the weak charge correction is currently in cesium PV experiments, where $\delta Q_{W}/Q_{W}\approx-0.8\%$ exceeds both the experimental error 0.35% and the error in our atomic many-body calculation ¡ 0.5%. This correction actually eliminates $2\sigma$ difference between the measured $Q_{W}^{\text{exp}}=-72.41(42)$ and the standard model $Q_{W}^{SM}+\delta Q_{W}^{SM}=-73.26(1-0.008)=-72.67$ weak charges.
After including $\delta Q_{W}^{SM}$, the measured in Cs PV experiment value of the Weinberg angle at $q^{2}\approx 0$,
$\sin^{2}\theta_{W}=0.2375(19)\,$, (7)
also comes to a perfect agreement with prediction of the standard model $\sin^{2}\theta_{W}=0.2387$.
The difference $Q_{W}^{\text{exp}}-(Q_{W}^{SM}+\delta Q_{W}^{SM})=0.26(42)$ allows us to obtain limits on new physics. The limits on the additional $Z^{\prime}$-boson interaction strength obtained as a function of $Z^{\prime}$ mass from the measurements of PV in Cs atom are approximately 2 times stronger than that presented in Table II and Fig. 1 of Ref. [13].
We also obtain limits on oblique radiative corrections produced by new particles. Ref. [31] calculated the shift of the weak charge produced by these corrections $\Delta Q_{W}=-0.8S$, where $S$ is the Peskin–Takeuchi parameter characterizing isospin-conserving vacuum-polarization effects. Comparing this expression with the measured value $\Delta Q_{W}=0.26(42)$, we obtain
S=0.32(53).S=-0.32(53)\,. (8)
The global electroweak fit of $S$ [16] is dominated by the $Z$-pole and $W$-mass measurements, while atomic PV probes oblique radiative corrections at very low momentum transfer, $q^{2}\approx 0$, thereby testing their running over a wide range of scales. Light weakly interacting states can induce momentum-dependent vacuum-polarization effects that contribute differently at $q^{2}\approx 0$ and $q^{2}\sim m_{Z}^{2}$. In this case, PV in cesium provides sensitivity to new physics effects that may not be fully captured by high-energy fits.
Acknowledgments. – The work was supported by the Australian Research Council Grant No. DP230101058.
Appendix A Reduction of effective interaction potentials to the contact interaction
The exchange of various SM particles other than photon between the atomic nucleus and the electrons may be described by potential forces which may be considered as a perturbation to the Coulomb interaction potential. If such a potential $V(r)$ is short-range as compared with the size of the nucleus, they may be effectively reduced to the contact interaction with the the delta-function potential,
$V_{\text{contact}}(\vec{r})=\kappa\delta^{3}(\vec{r})\,$, (9)
with some coefficient $\kappa$ defined by
$\kappa=\int V(r)d^{3}r\,.$ (10)
The canonical example is the exchange of the $Z$ boson which is described by the potential
$V_{Z}(r)=-\frac{g^{2}}{64\pi c_{W}^{2}}\frac{e^{-m_{Z}r}}{r}\,$, (11)
where $g\approx 0.65$ is the gauge coupling constant of the $SU(2)_{L}$ group, and $c_{W}\equiv\cos\theta_{W}\approx 0.87$ is the cos of Weinberg angle. Since the range of this interaction is $r=O(m_{Z}^{-1})$, in atomic calculations it may be replaced with the contact interaction (9) with the coefficient
$\kappa_{Z}=-\frac{g^{2}}{16c_{W}^{2}m_{Z}^{2}}\,.$ (12)
Below, we consider effective potentials corresponding the exchange of neutrinos as per the Feynman graphs in Fig. 1. In the present work, we need two such potentials, which are usually referred to as the self-energy (SE) and ‘penguin’ (PG) contributions, respectively. These potentials were found in the works [22, 23]. We will calculate the corresponding contact interaction coefficients $\kappa$ for these potentials.
A.1 Neutrino self-energy diagram
The potential corresponding to the neutrino one-loop self-energy diagram was calculated in Refs. [22, 23]:
$V_{\text{SE}}(r)=-V_{0,\text{SE}}\frac{1}{r}\int_{0}^{\infty}dt\frac{te^{-r\sqrt{t}}}{(t-m_{Z}^{2})^{2}}\,$, (13)
where
$V_{0,\text{SE}}=\frac{g^{4}}{(4c_{W})^{4}}\frac{1}{24\pi^{3}}\,.$ (14)
The integral in Eq. (13) is divergent because of the double pole in the denominator. This double pole is regularized by a small shift along the imaginary axis, and by retaining the principal value of the integral [23]. As a result, one finds
$V_{\text{SE}}(r)=$ $\frac{V_{0}}{2r}\Big[e^{m_{Z}r}(2+m_{Z}r)\mathrm{Ei}(-m_{Z}r)$
$+e^{-m_{Z}r}(2-m_{Z}r)\mathrm{Ei}(m_{Z}r)+2\Big],$ (15)
where
$\mathrm{Ei}(x)\equiv-\int_{-x}^{\infty}t^{-1}e^{-t}dt\,.$ (16)
The asymptotic behavior of the potential (15) is
$V(r)|_{r\gg m_{Z}^{-1}}=-\frac{g^{4}}{2\pi^{3}c_{W}^{4}r^{5}}\,.$ (17)
This potential plays significant role at the distance $r\sim m_{Z}^{-1}$, and decays rapidly at $r\gg m_{Z}^{-1}$. Hence, in atomic calculations, this potential may be replaced with the contact interaction potential (9) where the coefficient $\kappa$ is found to be
$\kappa_{\text{SE}}=4\pi\int_{0}^{\infty}r^{2}V(r)dr|_{m_{\nu}\to 0}=\frac{4\pi V_{0,\text{SE}}}{m_{Z}^{2}}\,.$ (18)
Here the following identity has been used:
$\int_{0}^{\infty}\left[e^{x}(2+x)\mathrm{Ei}(-x)+e^{-x}(2-x)\mathrm{Ei}(x)+2\right]xdx=2\,.$ (19)
Comparing Eq. (18) with Eq. (12) we find
$\frac{\kappa_{\text{SE}}}{\kappa_{Z}}=-\frac{g^{2}}{96\pi^{2}c_{W}^{2}}\approx-5.9\times 10^{-4}\,.$ (20)
A.2 Triangle PG diagram contribution to the potential
The contribution to the potential from the triangle PG diagrams was found in Refs. [22, 23] in the following form:
$V_{\text{PG}}(r)=-V_{0,\text{PG}}\frac{1}{r}\text{Re}I(r)\,$, (21a)
$I(r)=\int_{0}^{\infty}dt\,e^{-\sqrt{t}r}$
$\times\frac{t(2m_{W}^{2}+3t)-2(m_{W}^{2}+t)^{2}\ln\left(1+\frac{t}{m_{W}^{2}}\right)}{t^{2}(t-m_{Z}^{2}+i0)}\,$, (21b)
where
$V_{0,\text{PG}}=\frac{g^{4}}{2048\pi^{3}c_{W}^{2}}\,.$ (22)
Here the principal value of the integral is assumed near the pole $t=m_{Z}^{2}$. In Ref. [23], it was shown that the potential (21) possesses the following asymptotic behavior
$V_{\text{PG}}(r)|_{r\gg m_{Z}^{-1}}=-\frac{g^{4}}{256\pi^{3}c_{W}^{2}m_{Z}^{2}m_{W}^{2}r^{5}}\,.$ (23)
Thus, this potential drops rapidly and its effects become negligible at the distance of a few $m_{Z}^{-1}$. In atomic calculations, within the current accuracy goal the potential corresponding to the PG diagram may be replaced by the contact interaction (9) with the corresponding coefficient $\kappa$.
To compare the potential (21a) with contact interaction (9), we need to calculate the coefficient
$\kappa_{\text{PG}}=-4\pi V_{0,\text{PG}}\int_{0}^{\infty}I(r)r\,dr\,.$ (24)
Substituting here Eq. (21b) and integrating over $r$, we find
$I_{0}\equiv\int_{0}^{\infty}I(r)r\,dr$
$=\frac{1}{m_{W}^{2}}\int_{0}^{\infty}\frac{x(2+3x)-2(1+x)^{2}\ln(1+x)}{x^{3}(x-z)}\,.$ (25)
where $z\equiv\frac{m_{z}^{2}}{m_{W}^{2}}$. Making use of the identity
$\ln(1+x)=x^{2}-\frac{x^{2}}{2}+x^{3}\int_{0}^{1}\frac{\alpha^{2}d\alpha}{1+\alpha x}\,$, (26)
reduces the problem to the integration of rational functions,
$I_{0}=$ $\frac{1}{m_{W}^{2}}\int_{0}^{\infty}\frac{xdx}{x-z}$
$-\frac{2}{m_{W}^{2}}\int_{0}^{1}\alpha^{2}d\alpha\int_{0}^{\infty}\frac{(1+x)^{2}dx}{(1+\alpha x)(x-z)}\,.$ (27)
It is possible to show that all divergent terms in these integrals cancel, and the finite result is
$I_{0}=$ $\frac{1}{2z^{3}m_{W}^{2}}\Big[4(z+1)^{2}\text{Li}_{2}(-z)$
$+4(z+1)^{2}\ln(z)\ln(z+1)$
$+z(7z-2(3z+2)\ln(z)+4)\Big].$ (28)
Recalling that $z=c_{W}^{-2}$, we find
$I_{0}=$ $\frac{1}{2m_{Z}^{2}}\Big[4\left(c_{W}^{2}+1\right)^{2}\text{Li}_{2}\left(-c_{W}^{-2}\right)$
$+4c_{W}^{2}-2\log\left(c_{W}\right)\big(-4c_{W}^{2}$
$+4\left(c_{W}^{2}+1\right)^{2}\log\left(c_{W}^{-2}+1\right)-6\big)+7\Big]$
$\approx$ $-\frac{0.92}{m_{Z}^{2}}\,.$ (29)
As a result, the coefficient (24) is found to be:
$\kappa_{\text{PG}}\approx 15V_{0,\text{PG}}m_{Z}^{-2}\approx 5.4\times 10^{-5}m_{Z}^{-2}\,.$ (30)
Comparing this coefficient with Eq. (12), we find
$\frac{\kappa_{\text{PG}}}{\kappa_{Z}}\approx-1.6\times 10^{-3}\,.$ (31)